OpenStudy (anonymous):

here were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $400. How many sophomores bought tickets?

Parth (parthkohli):

Let's say that the number of freshmen is \(f\) and sophomores is \(s\).\[f + s = 150\\2f + 3s = 400 \]Solve the two equations.

Parth (parthkohli):

You do know how I got these equations, right?

OpenStudy (anonymous):

no :/

Parth (parthkohli):

All right. Do you understand how \(f + s = 150\)?

OpenStudy (anonymous):

because there were 150 freshman and sophomores that got tickets ?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

k, i got 100 is that right?

Parth (parthkohli):

100?

OpenStudy (anonymous):

yes

Parth (parthkohli):

100 for?

OpenStudy (anonymous):

100 sophomores and 50 for freshmen

Parth (parthkohli):

Yes that is correct!

OpenStudy (anonymous):

okay, thanks so much :)

Parth (parthkohli):

You're welcome =)